A medical investigation claims that the average number of infections per week

at a hospital is 16.3. A random sample of 10 weeks had a mean number of 17.7

infections. The sample standard deviation is 1.8. Is there enough evidence to

reject the investigator’s claim at α = 0.05? Assume the variable is normally

distributed.



1
Expert's answer
2021-06-24T07:24:33-0400

This is a one-sample t-test.

"H_0:\mu=16.3."

"H_a:\mu\ne 16.3."

Test statistic: "t=\frac{17.7-16.3}{\frac{1.8}{\sqrt{10}}}=2.46."

P-value: "p=2P(T>2.46)=0.0362."

Since the p-value is less than 0.05, reject the null hypothesis.

There is enough evidence to reject the investigator’s claim at α = 0.05.